OpenStudy (anonymous):

SOS Differential Equations! y'+y=xy^3

OpenStudy (anonymous):

\[y'+y=xy^3\]

OpenStudy (anonymous):

the answer is \[x=cy+\ln^2y\]

OpenStudy (anonymous):

it's clearly a Bernoulli equation

OpenStudy (chaise):

Looks like you answered your own question!

OpenStudy (anonymous):

@chaise It's from the book. but I am very far from the solution

OpenStudy (anonymous):

@IrishBoy123 ? Will you be my savior?

OpenStudy (anonymous):

@SithsAndGiggles ?

OpenStudy (anonymous):

Substitute \(t=y^{-2}\).

OpenStudy (anonymous):

t?

OpenStudy (anonymous):

you mean adter dividing by y^3?

OpenStudy (anonymous):

after

OpenStudy (anonymous):

Any letter or symbol will do. Dividing through by \(y^3\) yields \[y^{-3}y'+y^{-2}=x\] and substituting \(t=y^{-2}\) gives \(t'=-2y^{-3}y'\) via the chain rule, and so the ODE is linear in \(t\): \[-\frac{1}{2}t'+t=x\]

OpenStudy (anonymous):

oh yeah I did that

OpenStudy (anonymous):

and from there I'm quite confused tbh

OpenStudy (anonymous):

Well it's an ordinary linear equation, so you can solve it by finding an integrating factor, or undetermined coefficients, or some more nonstandard creative way.

OpenStudy (chaise):

It would be easier to multiply everything by -2, and then use integrating factor to solve: \[t'-2t=-2x\] Now apply integrating factor.

OpenStudy (chaise):

Let p(x)=e^int(-2.dt)=e^(-2t) Now apply reverse product rule. This video might help: https://www.youtube.com/watch?v=eWLgWazTc-0

OpenStudy (anonymous):

well I did went from bernoulli to a linear equation by the book, and it's very far away from the solution... can you solve it to see if you have the same problem?

OpenStudy (anonymous):

@SithsAndGiggles I did the same way, but I'm not sure if the problem is with me or with the answer

OpenStudy (anonymous):

@wio?

OpenStudy (anonymous):

Any idea?

OpenStudy (anonymous):

\[ \mu (x) = e^{\int p(x) dx} = e^{\int -2\;dx} = Ce^{-2x} \]

OpenStudy (anonymous):

Multiply by integrating factor: \[ \mu(x)t' + -2t\mu(x) = \mu(x)2x \implies (\mu(x)t)' = \mu(x) 2x \\ Ce^{-2x}t = \int Ce^{-2x}(-2x)\;dx \]Then integrate it.

OpenStudy (anonymous):

why did you go from \[Ce^-2x=\int\limits?\]

OpenStudy (anonymous):

\[ e^{\int -2\;dx} = e^{-2x+k} = Ce^{-2x} \]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Basically going backwards: \[ (Ce^{-2x}t) ' = (Ce^{-2x})'t + Ce^{-2x}t' = (Ce^{-2x})(-2t) +(Ce^{-2x})t' \]

OpenStudy (anonymous):

Then I integrate both sides. The integrating factor is designed this way.

OpenStudy (anonymous):

I'm a bit confused tbh, I don't see it leads me to the same answer as in the book..., I've just looked and I did just the same

OpenStudy (anonymous):

I guess then there's a problem with the answer in the book

OpenStudy (anonymous):

Multiply by integrating factor: \[ Ce^{-2x}t = \int Ce^{-2x}(-2x)\;dx \]Then integrate it. We get \[ Ce^{-2x}t = Ce^{-2x}\left(x+\frac{1}{2}\right) + K \implies t = x+\frac 12+ K'e^{2x} \]I'm guessing at this point we would undo our substitution.

OpenStudy (anonymous):

see? I did get that as well!!!! Horray!

OpenStudy (anonymous):

then it probably means the answer in the book is wrong

OpenStudy (anonymous):

thanks a lot! @wio !!!!

OpenStudy (irishboy123):

@Hipocampus well done, i go that too but i thought i'd stay out while you had all the big guns helping you out. those homogeneous eqns you were doing a few weeks back must seem real simple now!

OpenStudy (anonymous):

@IrishBoy123 thanks to you they are!!!!!!

OpenStudy (anonymous):

without your help I don't know what I would have done!

OpenStudy (anonymous):

wherever you are on the globe, thank YOU!